Please help.
Solve by addition.
2x-9y=-4
x-4y=-1

Answer by chitra(359)

The given equations are:

2x - 9y = -4 --------------------(1)
x - 4y = -1 --------------------(2)

Given that the solving should be done by addtion.

So lets multiply eqn(2) by (-2) and then add it with eqn (1)

2x - 9y = -4

-2x + 8y = 2
=============
0x - 1y = - 2

==> y = 2

And by back substitution, we get:


x - 4(2) = - 1

x = -1 + 8

x = 7

Hence, the solution.

Regards
==